A manufactuer of sports equip.Expects that 5 out of every 126basketballs manufactured will hav a defect. At this rate when 630 basketballs are produced how many will b defective?
Answer by stanbon(75887)

A manufactuer of sports equip.Expects that 5 out of every 126basketballs manufactured will hav a defect. At this rate when 630 basketballs are produced how many will b defective?
----
x/630 = 5/126
---
x = 630(5/126)
---
x = 25 (# of defective)
=====================
cheers,
Stan H.